OpenStudy (goformit100):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (1, 5) and (2, 9) (6, 2) and (3, 4) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

set #1 (1,5) and (2,9) set #2 (6,1) and (3,4) do you know how to calculate the slope of each pair?

OpenStudy (goformit100):

hOW ?

OpenStudy (anonymous):

you need to write the fraction \[slope_{PQ}=\frac{y_2-y_1}{x_2-x_1},where \space P(x_1,y_2),Q(x_2,y_2)\] using the two points given

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

i will do set #1 and you will do 2 so for set #1 the points are (1,5) and (2,9) so the change in y \[Δy=y_2−y_1=9−5=4\] \[Δx=x_2−x_1=2−1=1\] then we divide the two \[slope=4/1=4\]

OpenStudy (anonymous):

do you understand

OpenStudy (goformit100):

Which option is correct ?

OpenStudy (goformit100):

@Jonask Reply

OpenStudy (anonymous):

for set #2 slope=3/-3=-1 4>-1 so set #1> set#2